find the 37th term of the sequence 2,5,8,11,14
Answer by ankor@dixie-net.com(12704)

find the 37th term of the sequence 2,5,8,11,14
:
x = no of the term; y = value of the term
:
y = 3(x-1) + 2
y = 3x - 3 + 2
y = 3x -1
:
y = 3(37) - 1
y = 111 - 1
y = 110 is the 37th term